# Dynamic Attention Analysis for Backdoor Detection in Text-to-Image Diffusion Models

### v2

@@ III-D Dynamical System-based Approach
where a smaller $\gamma_{i}$ indicates lower stability. The term $A^{t}$ in (11) is the Laplacian matrix $A^{t}=[a^{t}_{ij}]_{i,j=1}^{L}$, where:
@@
It can be proven that $\frac{dV[X(t),t]}{dt}<0$ and the proof is provided in Appendix-B. Thus, the system is globally asymptotically stable, ensuring that the system stabilizes within a finite time [64]. To further validate this conclusion experimentally, we visualize the average derivative of the Lyapunov function over time for five backdoor scenarios and benign samples in Fig. 5. As can be seen, the derivative remains below zero over the time steps. Additionally, we observe that the absolute value of these derivative first increase and then decrease. It indicates that the diffusion generation process becomes increasingly chaotic in the early stages, e.g., 10 steps earlier, and then gradually converges to stability.
Since the state equation in (11) is a first-order ordinary differential equation (ODE), we numerically solve it to obtain the state function of each node at time step $t$. Specifically, we apply the RKF4(5) [65] method to solve it and obtain the dynamic state $x_{i}(t)$ for each node $v_{i}^{t}$. Inspired by the DAA-I, we refine the Evolve Rate (ER) of each map by:
@@
where $t$ is the start time step, with $s$ denoting the time span. By setting a threshold $\alpha_{G}\in\mathbb{R}^{1}$, we can distinguish whether a sample $P$ is a backdoor sample by:
[…]
We can observe that both (7) and (17) focus on the relative evolution rate of $<$EOS$>$ token. DAA-I individually computes the dynamic feature of each attention map, brings a simple but effective solution. On the other hand, DAA-S aims to fuse the dynamic correlations into features. Although it involves more complex computations, it provides a refined way to model global dynamic features of attention maps.
Figure 5: The average derivative of the Lyapunov function over time steps for five attack scenarios and benign samples. Each node represents the derivative at time step $t$.
@@ IV-A Experiment Settings
Attack configurations. Here, we consider five representative backdoor attack scenarios, including Rickrolling [23], Villan Diffusion [25], BadT2I [32], EvilEdit [28] and IBA [35]. In particular, for Rickrolling [23], we set the loss weight $\beta$ to 0.1 and fine-tune the text encoder for 100 epochs with a clean batch size of 64 and a backdoor batch size of 32. For Villan Diffusion [25], we fine-tune the model on CelebA-HQ-Dialog dataset [66] with LoRA [31] rank as 4. We train the model for 50 epochs with the training batch size as 1. For BadT2I [32], we inject backdoor to generate a specific object and fine-tune the model for 8000 steps with batch size as 8. For EvilEdit [28], we align 32 cross-attention layers in the UNet to inject backdoors. For IBA [35], we train the backdoor with a learning rate of 1e-4 for 600 epochs. The hyperparameters $\gamma$ and $\lambda$ of IBA are set to 1 and 0.01, respectively. Following the original settings [23, 33], we use the stable diffusion v1.4 (sd14) [5] as the victim model.
@@
where $d(\cdot,\cdot)$ represents the cosine similarity between feature embeddings. Only when the indicator functions of both CLIP and BLIP are equal to 1, the sample is included. Finally, we construct a dataset containing five backdoor attack scenarios with 6,500 backdoor samples and corresponding 6,500 benign samples.
@@
Method Rickrolling Villan EvilEdit IBA BadT2I Rickrolling Villan EvilEdit IBA BadT2I Avg F1 (%) ↑ Avg AUC (%) ↑
FTT [33] 96.94 80.00 7.75 0.71 22.22 99.98 88.91 57.06 94.12 59.86 41.52 79.99
@@
Baselines. We compare our method with the current state-of-the-art (SOTA) backdoor detection methods, including T2IShield-FTT [33], T2IShield-CDA [33], and UFID [34]. Specifically, for FTT, we set the threshold to $2.5$. For CDA, we use a pre-trained linear discriminant analysis model. For UFID, each sample produce 15 variations and compute the images’ similarity by CLIP [29]. We compute the feature distribution of 3,000 benign samples from DiffusionDB [68] and set the threshold of UFID to $0.776$.
Implementation details. We train both DAA-I and DAA-S on train set to find the best $t$, $n$, $\alpha_{I}$ and $\alpha_{G}$. For the DAA-I, we set $t=3$ and $s=2$. For the DAA-S, we set $t=1$ and $s=5$. The thresholds $\alpha_{I}$ and $\alpha_{G}$ are set to $0.00048$ and $0.00150$, respectively. Besides, the coupling strength $c$ is set to 1. The dynamical matrix $F=\{-1,-1,\dots,-10\}$. Denoising time steps $T$ for each sample is set to 50. The size of an attention map $D$ is $16$.
IV-B Detection Results
Table II shows the comparison of our proposed DAA-I and DAA-S with state-of-the-art (SOTA) methods in terms of F1 Score and AUC. First, we observe that baseline methods [33, 34] are effective in detecting classic backdoor attack methods, i.e., Rickrolling [23] and Villan [25]. However, their performance significantly degrades on the latest backdoor attacks, including EvilEdit [28], IBA [35], and BadT2I [32]. This decline is primarily due to the fact the static anomalies on which these detections rely are effectively mitigated by these attack methods. In contrast, our DAA yields competitive results in detecting classical backdoors and exhibits superior performance in newer attack scenarios. DAA-I achieves an average F1 Score of 74.54% and an average AUC of 86.17%. DAA-S achieves the best detection performance with an average F1 Score of 79.49% and an average AUC of 87.67%. Both methods significantly outperforming the previous state-of-the-art F1 Score of 41.52% and AUC of 79.99%, demonstrating the effectiveness by leveraging dynamic attention features for backdoor detection. Furthermore, the improvement observed with DAA-S highlights the importance of modeling the spatial correlations among attention maps in enhancing detection performance.
@@ IV-B Detection Results
Figure 8: Sensitivity to the base model in terms of (a) F1 Score (%) and (b) AUC (%). The blue bars represent results on SD14 [6], the yellow bars represent results on SD15 [6], and the red bars represent results on RV6 [69].
(a) DAA-I.
(b) DAA-S.
Figure 9: The generalization of DAA. Each number in the rows represents the score of AUC (%) for the corresponding attack scenario, with the last column showing the average AUC across five scenarios.
@@ IV-C Sensitivity to Generation Parameters
TABLE III: Efficiency analysis of detection methods. $U$ and $V$ represent the number of image generations and denoising time steps per sample, respectively. $T_{DI}$ represents the detection model inference time.
Method $U\times V$downarrow $T_{DI}$ (ms) $\downarrow$ All Time Cost (s) $\downarrow$
FTT [33] 1$\times$50 0.09 3.18
CDA [33] 1$\times$50 12.59 3.21
UFID [34] 15$\times$50 101.70 47.22
DAA-I 1$\times$5 0.27 0.32
DAA-S 1$\times$6 39.06 0.41
@@
t=0 74.72 76.62 79.15 81.23 82.49 82.62 82.72 82.50 82.23
t=1 42.03 80.75 81.94 82.68 82.76 82.19 81.68 81.33 81.19
t=2 80.08 81.91 82.74 82.77 82.13 81.59 81.48 81.29 81.05
DAA-I t=3 79.11 83.13 82.38 81.70 81.37 81.20 80.63 79.98 79.68
t=0 77.20 81.41 83.28 82.40 82.54 82.74 82.71 83.25 83.13
t=1 83.07 81.62 82.77 83.08 83.46 83.42 83.26 83.32 83.26
t=2 83.12 83.22 82.23 81.37 78.98 0.10 0.10 0.10 0.10
DAA-S t=3 80.85 44.35 0.03 0.03 0.03 0.00 0.35 0.52 0.56
Sensitivity to base model types. Given the plug-and-play nature of the backdoor-infected text encoder across various base models [23], we investigate the performance of DAA in detecting backdoor samples generated by different base models. In this experiment, we employ a text encoder injected with the Rickrolling trigger [23] and replace the original benign text encoders of SD14 [6], SD15 [6], and Realistic Vision V6 (RV6) [69]. We train the DAA on SD14 dataset and directly examine the performance of DAA on SD15 and RV6. As shown in Fig. 8, DAA demonstrates consistent detection performance across all tested base models. Specifically, DAA-I maintains detection performance above 80% for each model, while DAA-S consistently achieves detection performance exceeding 90%. The slight performance drop observed with RV6 may be due to extensive fine-tuning, which affects the anomalous features of the backdoor samples and lead to a decrease in detection performance. Nevertheless, the results indicate that the $<$EOS$>$ token’s anomalous behavior is a strong feature for detection, regardless of the specific base model. Furthermore, the superior performance of DAA-S suggests that it exhibits less sensitivity to base model types compared to DAA-I.
@@
DAA-S F1 Score (%) $\uparrow$ AUC (%) $\uparrow$ Avg F1 (%) $\uparrow$ Avg AUC (%) $\uparrow$
Rickrolling Villan EvilEdit IBA BadT2I Rickrolling Villan EvilEdit IBA BadT2I
$\gamma_{L}=-0.5$ 97.19 95.96 74.42 68.10 59.85 99.24 99.43 79.18 97.16 58.58 79.10 86.72
$\gamma_{L}=-1$ 96.94 96.79 73.44 68.57 61.15 98.93 99.46 79.33 96.83 59.15 79.38 86.74
[…]
$\gamma_{L}=-5$ 98.19 97.07 75.51 66.18 60.23 99.81 99.46 81.02 97.49 58.49 79.44 87.25
[…]
$\gamma_{L}=-15$ 98.19 95.11 74.42 69.50 62.56 99.43 99.47 80.06 96.95 60.56 79.96 87.29
@@
DAA-S F1 Score (%) $\uparrow$ AUC (%) $\uparrow$ Avg F1 (%) $\uparrow$ Avg AUC (%) $\uparrow$
Rickrolling Villan EvilEdit IBA BadT2I Rickrolling Villan EvilEdit IBA BadT2I
$c=0.5$ 96.94 93.66 72.44 73.65 62.46 97.92 99.48 81.42 96.34 58.63 79.83 86.76
$c=1$ 98.96 96.51 74.82 66.67 60.47 99.80 99.49 90.19 97.54 61.26 79.49 87.67
$c=5$ 97.94 95.11 70.75 80.26 65.35 99.89 99.28 74.15 98.23 59.57 81.88 86.22
[…]
@@
DAA-S 87.67 64.34 68.20
IV-D Efficiency
In this section, we compare the efficiency of DAA with several baseline methods. We conduct the experiments on a RTX4090-24GB GPU and the detailed implementation conditions are provided in Appendix-A. We apply different detection methods on 100 samples and calculate the average time cost per sample. For a backdoor detection method on the text-to-image diffusion model, the all time cost per sample is proportional to:
$All\ Time\ Cost\propto(U\times V\times T_{DE}+T_{DI}),$ (20)
For T2IShield, it requires $U=1$ generation and additional detection inference time for the “Assimilation Phenomenon.” For UFID, it takes $U=15$ generations and additional detection inference time to compute graph diversity. In contrast, since DAA only focuses on the early stages in denoising process, the detection is completed before the image generation. We present the detailed time cost in Table III. As can be seen, FTT achieves the fastest speed on $T_{DI}$, only requiring 0.09 ms for calculating the feature. DAA-I takes 0.27 ms to model the dynamic feature of the attention maps, while DAA-S takes 39.06 ms. Although DAA requires more time cost on detection inference, since the $T_{DE}\gg T_{DI}$, DAA achieve a significant low time consumption compared to previous methods. DAA-I requires about 10% time cost of FTT and DAA-S requires about 13%. These results highlight the efficiency advantage of DAA over existing methods, making it a practical solution for real-time backdoor detection in text-to-image diffusion models.
IV-E Generalization
To study the generalization of the proposed methods, we train the models on one backdoor scenario and test them on others. Fig. 9 shows the generalization of DAA-I and DAA-S respectively. Each number in the rows represents the AUC value for the corresponding attack scenario, with the last column showing the average AUC across the five scenarios. Darker colors indicate higher values. As observed, the two methods exhibit different generalization capabilities. For DAA-I, the differences across each column are not significant, suggesting that the detection performance of this method is less sensitive to the distribution of the training data. We also notice that the detection performance on BadT2I is poor across all rows, which suggests that BadT2I is a challenging backdoor attack scenario for DAA-I.
Additionally, we observe that DAA-S achieves the best generalization performance when trained on the Rickrolling [23] and Villan [25] attack scenarios, with average AUC scores of 90.26% and 90.27%, respectively. However, in the remaining three scenarios, DAA-S exhibits poorer generalization performance compared to DAA-I. This highlights the variability in DAA-S’s performance across different attack types, suggesting that while it excels in certain contexts, its generalization capability is not as robust as that of DAA-I.

### v3

@@ III-D Dynamical System-based Approach
where a smaller $\gamma_{i}$ indicates lower stability. The term $A^{t}$ in Eq. (11) is the Laplacian matrix $A^{t}=[a^{t}_{ij}]_{i,j=1}^{L}$, where:
@@
It can be proven that $\frac{dV[X(t),t]}{dt}<0$ and the proof is provided in Appendix-B. Thus, the system is globally asymptotically stable, ensuring that the system stabilizes within a finite time [70]. To further validate this conclusion experimentally, we visualize the average derivative of the Lyapunov function over time for six backdoor scenarios and benign samples in Fig. 5. As can be seen, the derivative remains below zero over the time steps. Additionally, we observe that the absolute value of these derivative first increase and then decrease. It indicates that the diffusion generation process becomes increasingly chaotic in the early stages, e.g., 10 steps earlier, and then gradually converges to stability.
Since the state equation in Eq. (11) is a first-order ordinary differential equation (ODE), we numerically solve it to obtain the state function of each node at time step $t$. Specifically, we apply the RKF4(5) [71] method to solve it and obtain the dynamic state $x_{i}(t)$ for each node $v_{i}^{t}$. Inspired by the DAA-I, we refine the Evolve Rate (ER) of each map by:
@@
Figure 5: The average derivative of the Lyapunov function over time steps for six attack scenarios and benign samples. Each node represents the derivative at time step $t$.
@@
where $t$ is the start time step, with $s$ denoting the time span. By setting a threshold $\alpha_{S}\in\mathbb{R}^{1}$, we can distinguish whether a sample $P$ is a backdoor sample by:
$\text{DAA-S}(P)=\mathbb{I}[\text{DAA-S}^{t}_{s}\leqslant\alpha_{S}].$ (18)
We further visualize the distribution of $\text{DAA-S}^{1}_{5}$. As shown in Fig. 4b, similarly, we also observe a distribution shift between backdoor and benign samples. By comparing the distribution results between Fig. 4a and Fig. 4b, we find that BadT2I [32] and EvilEdit [28] are the two backdoor samples that exhibit dynamic characteristics more similar to benign samples. Besides, we observe that under the DAA-S method, features for each method show a distribution that is closer to a normal distribution, whereas the DAA-I exhibits skewed distribution. This suggests that DAA-S is a more effective approach for modeling dynamic anomaly.
We can observe that both Eq. (7) and Eq. (17) focus on the relative evolution rate of <<EOS>> token. DAA-I individually computes the dynamic feature of each attention map, brings a simple but effective solution. On the other hand, DAA-S aims to fuse the dynamic correlations into features. Although it involves more complex computations, it provides a refined way to model global dynamic features of attention maps. Compared with the vector-based formulation in Eq. (5), both methods compute a scalar metric for each sample and determine it using a statistically estimated threshold. This design enables more efficient backdoor sample detection and facilitates deployment in real-world scenarios.
@@ IV-A Experiment Settings
Attack configurations. Here, we consider six representative backdoor attack scenarios, including Rickrolling [23], Villan Diffusion [25], BadT2I [32], EvilEdit [28], IBA [35] and PersonBA [24]. In particular, for Rickrolling [23], we set the loss weight $\beta$ to 0.1 and fine-tune the text encoder for 100 epochs with a clean batch size of 64 and a backdoor batch size of 32. For Villan Diffusion [25], we fine-tune the model on CelebA-HQ-Dialog dataset [72] with LoRA [31] rank as 4. We train the model for 50 epochs with the training batch size as 1. For BadT2I [32], we inject backdoor to generate a specific object and fine-tune the model for 8000 steps with batch size as 8. For EvilEdit [28], we align 32 cross-attention layers in the UNet to inject backdoors. For IBA [35], we train the backdoor with a learning rate of 1e-4 for 600 epochs. The hyperparameters $\gamma$ and $\lambda$ of IBA are set to 1 and 0.01, respectively. For PersonBA [24], we adopt a learning rate of 5e-04, with 2,000 training steps and a batch size of 4. Each backdoor is constructed using 4 images to represent a specific object. Following the original settings [23, 33], we use the stable diffusion v1.4 (sd14) [5] as the victim model.
@@
where $d(\cdot,\cdot)$ represents the cosine similarity between feature embeddings. Only when the indicator functions of both CLIP and BLIP are equal to 1, the sample is included.
Baselines. We compare our method with the current state-of-the-art (SOTA) backdoor detection methods, including T2IShield-FTT [33], T2IShield-CDA [33], UFID [34] and DisDet [63]. Specifically, for FTT, we set the threshold to $2.5$. For CDA, we use a pre-trained linear discriminant analysis model. For UFID, each sample produce 15 variations and compute the images’ similarity by CLIP [29]. We compute the feature distribution of 3,000 benign samples from DiffusionDB [74] and set the threshold of UFID to $0.776$. For DisDet, we compute the the distribution discrepancy of the noise input of 3,000 benign samples from DiffusionDB and set the threshold of DisDet to 0.00176.
@@
Method Rickrolling Villan EvilEdit IBA BadT2I PersonBA Rickrolling Villan EvilEdit IBA BadT2I PersonBA
Avg
F1 (%) ↑
Avg
AUC (%) ↑
DisDet [63] 0.0 88.57 0.0 0.0 5.83 0.0 51.57 93.51 54.52 50.42 53.35 70.30 15.73 62.28
FTT [33] 96.94 80.00 7.75 0.71 22.22 19.01 99.98 88.91 57.06 94.12 59.86 55.71 37.77 75.94
@@
Implementation details. We train both DAA-I and DAA-S on train set to find the best $t$, $n$, $\alpha_{I}$ and $\alpha_{S}$. For the DAA-I, we set $t=3$ and $s=2$. For the DAA-S, we set $t=1$ and $s=5$. The thresholds $\alpha_{I}$ and $\alpha_{S}$ are set to 0.0011-0.0011 and 0.0053-0.0053, respectively. Besides, the coupling strength $c$ is set to 1. The dynamical matrix $F=\{-1,-1,\dots,-10\}$. Denoising time steps $T$ for each sample is set to 50. The size of an attention map $D$ is $16\times 16$.
@@
Figure 8: Sensitivity to the base model in terms of (a) F1 Score (%) and (b) AUC (%). The bars represent results on SD14 [6], SD15 [6], RV6 [75], SDXL [76] and SD3 [77].
IV-B Detection Results
Table II shows the comparison of our proposed DAA-I and DAA-S with state-of-the-art (SOTA) methods in terms of F1 Score and AUC. First, we observe that baseline methods [33, 34] are effective in detecting classic backdoor attack methods, i.e., Rickrolling [23] and Villan [25]. However, their performance significantly degrades on the latest backdoor attacks, including EvilEdit [28], IBA [35], BadT2I [32] and PersonBA [24]. This decline is primarily due to the fact the static anomalies on which these detections rely are effectively mitigated by these attack methods. Besides, while DisDet [63] achieve nearly 100% accuary in detecting noise-like backdoor, it suffers a significant performance drop when facing text-based triggers in text-to-image backdoor attacks, achieving only average F1 Score of 15.73% and average AUC of 62.28%. This is because the textual trigger does not significantly alter the noise distribution, causing this detection method to fail in detecting backdoor samples in text-to-image diffusion models. In contrast, our DAA yields competitive results in detecting classical backdoors and exhibits superior performance in newer attack scenarios. DAA-I achieves an average F1 Score of 73.43% and an average AUC of 84.14%. DAA-S achieves the best detection performance with an average F1 Score of 79.27% and an average AUC of 86.27%. Both methods significantly outperforming the previous state-of-the-art F1 Score of 37.77% and AUC of 77.02%, demonstrating the effectiveness by leveraging dynamic attention features for backdoor detection. Furthermore, the improvement observed with DAA-S underscores the role of spatial correlations among attention maps in enhancing detection performance.
@@ IV-C Sensitivity to Generation Parameters
Sensitivity to base model types. Given the plug-and-play nature of the backdoor-infected text encoder across various base models [23], we investigate the performance of DAA in detecting backdoor samples generated by different base models. In this experiment, we employ a text encoder injected with the Rickrolling trigger [23] and replace the original benign text encoders of SD14 [6], SD15 [6], and Realistic Vision V6 (RV6) [75]. We also test a larger-scale model, i.e., SDXL [76] as well as a DiT-based model, i.e., SD3 [77]. We use the parameters trained on SD14 and directly apply them to other backbone models. As shown in Fig. 8, DAA achieves more than 72% in terms of F1 score in backdoor detection across all models. On SD15, the F1 results are consistent with those on SD14, achieving a F1 Score of 95.24% for DAA-I and 99.90% for DAA-S. We also observe some performance degradation on other models. This may be because the feature distributions of samples vary across different models, making the original thresholds less effective. Nevertheless, we find that DAA-S consistently maintained AUC values above 86% across all models, indicating that it can still effectively capture the discriminative features between backdoor and benign samples. The results indicate that the <<EOS>> token’s anomalous behavior is a strong feature for detection, regardless of the specific base model. Furthermore, the superior performance of DAA-S suggests that it exhibits less sensitivity to base model types compared to DAA-I.
TABLE III: Efficiency analysis of detection methods. $U$ and $V$ represent the number of image generations and denoising time steps per sample, respectively. $T_{DI}$ represents the detection model inference time. $T_{All}$ is the whole time cost per detection.
Method $U\times V$ $\downarrow$ $T_{DI}$ (ms) $\downarrow$ $T_{All}$ (s) $\downarrow$ Avg. AUC (%) $\uparrow$
DisDet [63] $\times1$ 0.27 0.06 62.28
[…]
IV-D Efficiency
In this section, we compare the efficiency of DAA with several baseline methods. We conduct the experiments on a RTX4090-24GB GPU and the detailed implementation conditions are provided in Appendix-A. We apply different detection methods on 100 samples and calculate the average time cost per sample. For a backdoor detection method on the text-to-image diffusion model, the all time cost $T_{All}$ per sample is proportional to:
$T_{All}\propto(U\times V\times T_{DE}+T_{DI}),$ (20)
For T2IShield, it requires $U=1$ generation and additional detection inference time for the “Assimilation Phenomenon.” For UFID, it takes $U=15$ generations and additional detection inference time to compute graph diversity. For Disdet, since it focuses on the distribution discrepancy of the denoising noise, it only requires $V=1$. Similarly, DAA only also focuses on the early stages in denoising process, the detection is completed before the image generation.
We present the detailed time cost in Table III. As can be seen, DisDet achieves the fastest speed on $T_{DI}$, only requiring 0.27ms for calculating the feature and taking 0.06s to complete the whole detection. However, the average AUC of Disdet only achieves 62.28%, making it fail to detect backdoor samples. DAA-I takes 0.27 ms to model the dynamic feature of the attention maps, while DAA-S takes 39.06 ms. Although DAA requires more time cost on detection inference, since the $T_{DE}\gg T_{DI}$, DAA achieve a significant low time consumption while the best detection performance compared to previous methods. DAA-I requires about 10% time cost of FTT and DAA-S requires about 13%. These results highlight the efficiency advantage of DAA over existing methods, making it a practical solution for effective and efficient backdoor detection in T2I diffusion models.
(a) DAA-I.
(b) DAA-S.
Figure 9: The generalization of DAA. Each number in the rows represents the score of AUC (%) for the corresponding attack scenario, with the last column showing the average AUC across six scenarios.
@@ IV-D Efficiency
t=0 69.72 71.49 73.12 75.78 76.50 76.52 76.37 76.13 76.09
t=1 66.66 74.47 76.19 76.12 76.05 76.38 75.95 75.39 75.13
t=2 74.53 76.24 75.79 76.10 76.33 75.75 75.84 75.83 75.11
DAA-I t=3 73.28 76.55 76.37 75.97 75.57 75.70 74.84 74.88 73.70
t=0 71.87 76.95 75.47 76.57 76.81 77.45 77.87 77.96 78.37
t=1 77.26 76.83 77.24 78.34 79.72 78.86 79.32 79.63 79.71
t=2 77.80 78.74 78.59 78.00 4.67 6.26 7.39 8.35 7.93
DAA-S t=3 78.00 4.43 4.47 5.23 8.38 8.15 8.49 8.09 10.39
IV-E Generalization
To study the generalization of the proposed methods, we train the models on one backdoor scenario and test them on others. Fig. 9 shows the generalization of DAA-I and DAA-S respectively. Each number in the rows represents the AUC value for the corresponding attack scenario, with the last column showing the average AUC across the six scenarios. Darker colors indicate higher values. As observed, the two methods exhibit different generalization capabilities. For DAA-I, the differences across each column are not significant, suggesting that the detection performance of this method is less sensitive to the distribution of the training data. We also notice that the detection performance on BadT2I is poor across all rows, which suggests that BadT2I is a challenging backdoor attack scenario for DAA-I.
Additionally, we observe that DAA-S achieves the best generalization performance when trained on the Rickrolling [23], Villan [25] and PersonBA [24] attack scenarios, with average AUC scores of 87.13%, 86.96% and 86.51%, respectively. However, in the remaining three scenarios, DAA-S exhibits poorer generalization performance compared to DAA-I. This highlights the variability in DAA-S’s performance across different attack types, suggesting that while it excels in certain contexts, its generalization capability is not as robust as that of DAA-I.
@@ IV-E Generalization
DAA-S F1 Score (%) $\uparrow$ AUC (%) $\uparrow$
Avg
F1 (%) $\uparrow$
Avg
AUC (%) $\uparrow$
Rickrolling Villan EvilEdit IBA BadT2I PersonBA Rickrolling Villan EvilEdit IBA BadT2I PersonBA
$\gamma_{L}=-0.5$ 96.94 93.66 74.10 66.67 60.22 71.34 97.96 99.38 82.45 96.29 58.54 87.88 77.16 87.08
$\gamma_{L}=-1$ 96.94 95.68 73.68 67.15 61.54 74.42 98.17 99.48 82.15 96.60 57.95 87.32 78.24 86.95
$\gamma_{L}=-2$ 97.19 95.68 73.68 69.04 64.13 73.21 97.47 99.44 81.48 96.45 56.06 88.03 78.82 86.49
$\gamma_{L}=-5$ 97.44 89.71 71.92 76.32 64.13 70.54 97.23 98.98 82.93 95.52 58.58 89.29 78.18 87.09
$\gamma_{L}=-10$ 97.44 98.95 73.68 72.03 62.82 70.72 99.80 99.49 80.19 97.54 61.26 79.35 79.27 86.27
$\gamma_{L}=-15$ 96.20 97.10 69.91 88.40 65.32 70.11 97.06 99.14 78.26 85.74 53.98 85.77 81.17 83.33
@@
DAA-S F1 Score (%) $\uparrow$ AUC (%) $\uparrow$
Avg
F1 (%) $\uparrow$
Avg
AUC (%) $\uparrow$
Rickrolling Villan EvilEdit IBA BadT2I PersonBA Rickrolling Villan EvilEdit IBA BadT2I PersonBA
$c=0.5$ 95.24 97.63 70.81 78.37 63.92 72.69 97.22 99.43 81.27 95.75 57.71 82.03 79.78 85.40
$c=1$ 97.44 98.95 73.68 72.03 62.82 70.72 99.80 99.49 80.19 97.54 61.26 79.35 79.27 86.27
$c=5$ 96.18 96.08 71.44 82.69 62.92 71.69 98.76 99.42 79.84 97.80 55.42 80.56 80.17 85.13
$c=10$ 97.19 96.79 73.38 78.18 63.69 71.98 98.31 99.46 81.12 97.78 55.94 82.49 80.20 85.68
@@
DAA-S 86.27 64.79 70.15
DAA-S F1 Score (%) $\uparrow$ AUC (%) $\uparrow$
Avg
F1 (%) $\uparrow$
Avg
AUC (%) $\uparrow$
Rickrolling Villan EvilEdit IBA BadT2I PersonBA Rickrolling Villan EvilEdit IBA BadT2I PersonBA
SSIM 85.45 98.95 64.84 26.76 26.75 70.59 90.72 99.83 82.05 84.31 70.01 74.54 62.22 83.58
1-norm 92.01 96.60 70.34 60.08 65.67 70.93 96.34 99.36 80.99 72.85 54.97 87.53 75.94 82.01
F-norm 97.44 98.95 73.68 72.03 62.82 70.72 99.80 99.49 80.19 97.54 61.26 79.35 79.27 86.27